OpenStudy (anonymous):

A bag of pins has 4 black, 2 brown and 6 blue pins in it. If 3 pins are selected without replacement then determine the probability of getting black, brown and blue in that order?

OpenStudy (anonymous):

please help me i would really appreciate it :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the number of ways to select pins such that one with each color is=4C1*2C1*6C1 the number of ways to select three =12C3 the probability is 4C1*2C1*6C1/12C3

OpenStudy (anonymous):

i don't have calculator so u might solve it

OpenStudy (anonymous):

yourself

OpenStudy (anonymous):

thank you for the help but i dont really understand what the C meas

OpenStudy (anonymous):

combination=C there is a buton on ur calculator

OpenStudy (anonymous):

ah ok thank you im not sure if i have that cuz i live in australia but im sure i will work it out once again thank you so much for helping

OpenStudy (anonymous):

plz give me medel if i helped i really need a bit of those to show them to my girlfriend

OpenStudy (anonymous):

ok sure thing how do that.. sorry new to this

OpenStudy (anonymous):

let me tell u

OpenStudy (anonymous):

o thanks

OpenStudy (anonymous):

did i do it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh cool

OpenStudy (raden):

hmmmm... in sentences told, the color are black, brown and blue. but ur question, the color are red, green and blue. not sincron :p

OpenStudy (anonymous):

oh sorry i was just wanted to change the colours cuz i dont like red or green i know im weird

OpenStudy (anonymous):

also what does sincron mean

OpenStudy (raden):

nopes... nevermind :) so, whats problem now

OpenStudy (raden):

is even take pins one by one ?

OpenStudy (raden):

if yes, it should be P(1Black and 1brown and 1blue) = 4/12 * 2/11 * 6/10

OpenStudy (anonymous):

ok thank you and yeah i meant yes does * mean times

OpenStudy (raden):

it can be 48/1320 = 2/55

OpenStudy (anonymous):

ok thank you so much :)

OpenStudy (raden):

welcome